OpenStudy (kayluh123):

Simplify. -3(u+x)+5(-4x-5u)

OpenStudy (mathlegend):

Distribute that -3

OpenStudy (mathlegend):

I will do one set of distribution... you will do the other one. @kayluh123

OpenStudy (mathlegend):

-3(u+x) -3(u) = -3u -3(x) = -3x -3u - 3x Do you understand? All I did was distribute that negative 3

OpenStudy (kayluh123):

so is it the same for the second part too?

OpenStudy (kayluh123):

-20x-25u?

OpenStudy (mathlegend):

Yes!

OpenStudy (mathlegend):

-3u - 3x - 20x - 25u = Just combine like terms... so the x's together and the u's together... simplify this

OpenStudy (kayluh123):

-28u-23x?

OpenStudy (mathlegend):

Yes, good job!

OpenStudy (kayluh123):

thanks so much